You are considering buying a delivery van for your company. It will cost​ $35,000 right now​ (Year 0), and you estimate that with the​ van, your profits will be​ $7,500 per year for Years 1 through 10. At the end of Year​ 10, you think you can sell the van for​ $1,000. Assume a MARR of​ 18%.

a. What is the present worth of the​ van?

b. Is this investment worth​ considering?

a. What is the present worth of this​ equipment?

          

​$nothing  

​(Round to the nearest dollar. Note that this value could be positive or​ negative.)

b. Is this investment worth​ considering?

A.

No

B.

Yes
Answer #1

Below is the calculation of the PW of the van:

Year CF Discount Factor Discounted CF
0 $   -35,000.00 1/(1+0.18)^0= 1 1*-35000= -35,000.00
1 $       7,500.00 1/(1+0.18)^1= 0.847457627 0.847457627118644*7500=      6,355.93
2 $       7,500.00 1/(1+0.18)^2= 0.71818443 0.718184429761563*7500=      5,386.38
2 $       7,500.00 1/(1+0.18)^2= 0.71818443 0.718184429761563*7500=      5,386.38
4 $       7,500.00 1/(1+0.18)^4= 0.515788875 0.515788875151941*7500=      3,868.42
5 $       7,500.00 1/(1+0.18)^5= 0.437109216 0.437109216230459*7500=      3,278.32
6 $       7,500.00 1/(1+0.18)^6= 0.370431539 0.370431539178355*7500=      2,778.24
7 $       7,500.00 1/(1+0.18)^7= 0.313925033 0.313925033201996*7500=      2,354.44
8 $       7,500.00 1/(1+0.18)^8= 0.266038164 0.266038163730505*7500=      1,995.29
9 $       7,500.00 1/(1+0.18)^9= 0.225456071 0.225456070958055*7500=      1,690.92
10 $       7,500.00 1/(1+0.18)^10= 0.191064467 0.191064466913606*7500=      1,432.98
Present Worth = Sum of all Discounted CF        -472.70

As present worth is negative, the investment is not worth it, because the profits are not recovering the costs on a discounted cf basis at the MARR of 18%